Break-Even for a Service Firm

Jonah Graham owns and operates The Green Thumb Company (GTC), which provides live plants and flower arrangements to professional offices. Jonah has fixed costs of $3,800 per month for office/greenhouse rent, advertising, and a delivery van. Variable costs for the plants, fertilizer, pots, and other supplies average $22 per job. GTC charges $60 per month for the average job.

Required:

1. How many jobs must GTC average each month to break even?
jobs per month

2. What is the operating income for GTC in a month with 97 jobs? Enter a net loss as a negative amount.
$

What is the operating income for GTC in a month with 107 jobs?
$

3. Jonah faces a tax rate equal to 25 percent. How many jobs must Jonah have per month to earn an after-tax income of $1,380? Round your answer to the nearest whole number of jobs.
jobs per month

4. Suppose that Jonah’s fixed costs increase to $3,966 per month and he decides to increase the price to $76 per job. What is the new break-even point in number of jobs per month? Round your answer to the nearest whole number of jobs.
jobs per month
Answer #1
Required 1 :
Contribution margin per job = Charges per job - Variable cost per job = 60 - 22 38
Break-even point = Fixed costs per month / Contribution margin per job = 3800 / 38 100 jobs per month
Required 2 :
97 jobs
Sales ( 97 * 60 ) 5820
(-) Variable costs ( 97 * 22 ) 2134
Contribution margin 3686
(-) Fixed costs 3800
Operating loss -114
107 jobs
Sales ( 107 * 60 ) 6420
(-) Variable costs ( 107 * 22 ) 2354
Contribution margin 4066
(-) Fixed costs 3800
Operating income 266
Required 3 :
Before tax income required = After tax income required / ( 1 - Tax%) = 1380 / ( 1 - 25% ) 1840
Per month Jobs required = ( Before tax income required + Fixed costs per month ) / Contribution margin per job = ( 1840 + 3800 ) / 38 148
Required 4 :
Contribution margin per job = Charges per job - Variable cost per job = 76 - 22 54 per job
Break-even point = Fixed costs per month / Contribution margin per job = 3966 / 54    73 Jobs per month